OpenStudy (anonymous):

Help please?! There are 162 games in a major league baseball season. What percentage of the remaining 90 games would the Tigers have to win to finish with a .600 winning percentage? (Hint: This problem is a little different. The sum of 29 and some portion of their remaining games divided by 162 games has to equal .6.)

OpenStudy (anonymous):

Help would be greatly appreciated! :'(

OpenStudy (anonymous):

Does the question indicate that they have won 29 of their first 72 games?

OpenStudy (anonymous):

... Not quite sure. Would you be able to solve it if you went by that? ... Ugh. Sorry I'm not good help :/

OpenStudy (anonymous):

I think they did win 29 of their first 72 games. So 162 * .6 = 97.2. So that means 97.2 - 29 = 68.2. So rounding gives you 68.

OpenStudy (anonymous):

Solve the following for w. w is the percentage of games that have to be won in the remaining 90 games.\[\frac{290+9 w}{1620}=0.6 \]

OpenStudy (anonymous):

Thanks guys :)